# CallNavi: A Study and Challenge on Function Calling Routing and Invocation in Large Language Models

### v1

@@
Abstract
Interacting with a software system via a chatbot can be challenging, especially when the chatbot needs to generate API calls, in the right order and with the right parameters, to communicate with the system. API calling in chatbot systems poses significant challenges, particularly in complex, multi-step tasks requiring accurate API selection and execution. We contribute to this domain in three ways: first, by introducing a novel dataset designed to assess models on API function selection, parameter generation, and nested API calls; second, by benchmarking state-of-the-art language models across varying levels of complexity to evaluate their performance in API function generation and parameter accuracy; and third, by proposing an enhanced API routing method that combines general-purpose large language models for API selection with fine-tuned models for parameter generation and some prompt engineering approach. These approaches lead to substantial improvements in handling complex API tasks, offering practical advancements for real-world API-driven chatbot systems.
CallNavi: A Study and Challenge on Function Calling Routing and Invocation in Large Language Models
Yewei Song1, Cedric Lothritz2, Xunzhu Tang1, Saad Ezzini3, Jacques Klein1, Tegawendé F. Bissyandé1, Andrey Boytsov4, Ulrick Ble 4, Anne Goujon4, 1University of Luxembourg, 2Luxembourg Institute of Science and Technology, 3Lancaster University, 4BGL BNP PARIBAS, Correspondence: yewei.song@uni.lu
1 Introduction
@@
Modern conversational AI systems, such as chatbots, rely heavily on accurate API calling to enable effective user interactions. In complex scenarios, models must go beyond generating simple API calls to handle challenges such as selecting the correct API from an extensive list, orchestrating multiple sequential API calls, and managing nested calls. Despite significant progress in generating syntactically correct single API calls, there has been limited focus on tasks requiring a sequence of API calls with logical dependencies, a crucial capability for real-world applications.
@@
Large Language Models (LLMs) such as GPT-4 Achiam et al. (2023) and Llama Touvron et al. (2023) have demonstrated remarkable capabilities across a wide range of natural language processing tasks. By leveraging extensive pretraining on diverse datasets, LLMs can generate coherent and contextually relevant responses, making them essential components of modern conversational AI systems. However, while their abilities in tasks like summarization and question answering are well-established, their proficiency in generating structured outputs—such as API calls, program code, and other machine-readable formats—remains a challenging frontier.
@@
Structured output generation requires LLMs to produce outputs that adhere to predefined syntactic and semantic rules, a task significantly more constrained than generating free-form text. Recent works have explored structured output generation for applications like code generation Chen et al. (2021), table completion Herzig et al. (2020), and multi-turn dialogue Budzianowski et al. (2018). Tools like CodeX Chen et al. (2021) and AlphaCode Li et al. (2022) have specifically focused on structured programming tasks, demonstrating the potential of LLMs in generating functionally valid code. Similarly, methods such as chain-of-thought (CoT) prompting Wei et al. (2022) and tool-augmented reasoning frameworks Qin et al. (2023) have enabled LLMs to reason step by step, improving their accuracy in complex tasks.
@@
Several earlier works have explored function calling and tool-use capabilities in LLMs, laying a foundation for modern benchmarks. Early studies, such as BotBase Zamanirad et al. (2017), focused on translating natural language into API calls, demonstrating initial progress in automating tool use. More recent benchmarks, including API-Bank Li et al. (2023), ToolEyes Ye et al. (2024), and BFCL Yan et al. (2024), evaluate LLMs on their ability to execute API calls. For instance, API-Bank assesses tool-augmented models on real-world tasks but operates with fewer than five API candidates per task, limiting scenario complexity. Similarly, ToolBench Liu et al. (2024) and ToolEyes test multi-tool scenarios but lack support for nested or highly interdependent API calls.
@@
To address these gaps, we propose CallNavi, a novel benchmark designed to evaluate large language models (LLMs) on tasks involving:
  • API selection from an unfiltered list of over 100 candidates;
  • Execution of multiple sequential API calls;
  • Handling of nested API interactions.
@@
Unlike prior datasets, CallNavi introduces real-world complexity by simulating unfiltered API selection and combining generation with routing tasks. It categorizes questions into easy, medium, and hard levels, enabling the evaluation of models across varying levels of task complexity. Our benchmark also proposes new metrics, including stability score, to measure consistency in model predictions over multiple runs.
@@
We comprehensively benchmark 17 LLMs on CallNavi, covering diverse commercial, general-purpose, and fine-tuned models. Our findings reveal key insights into the strengths and limitations of current models, providing a foundation for further advancements in API selection and function calling.
2 Related Work
2.1 Function Calling and API Selection
Table 1 compares CallNavi with existing benchmarks across key dimensions, highlighting its unique contributions. Notably, our CallNavi is the only benchmark to test models on unfiltered API selection with over 100 candidates. And CallNavi incorporate both multi-call and nested API tasks in a single dataset, provide a diverse set of 10 domains while maintaining structured difficulty levels.
@@ 2.1 Function Calling and API Selection
Table 1: Comparison of CallNavi with existing API function-calling benchmarks test set.
@@
2.2 Techniques for Structured Output and Stability
Recent advances in structured output generation, such as MetaGPT and CodeAgent, have inspired innovations in generating API call sequences Hong et al. (2023); Tang et al. (2024). These works emphasize task decomposition and step-by-step reasoning, which align with our use of function-calling routing for complex API tasks.
Ensuring stability across multiple runs remains a critical challenge for function-calling tasks. Prior works addressing LLM stability Du et al. (2023); Tian et al. (2023) have informed our development of the stability score, which quantifies consistency in predictions. This metric complements traditional accuracy metrics like AST match and exact match, enabling a more comprehensive evaluation of model reliability in real-world scenarios.
Figure 1: Pipeline of API calling while using LLMs
3 Research Questions
@@ 4 CallNavi Dataset
All generated API functions were manually reviewed for accuracy, consistency, and relevance. Specifically:
@@
  • Coverage of straightforward, single-function usage scenarios.
@@
  • Hard Questions: Designed to address edge cases, ambiguous queries, and nested API calls requiring advanced reasoning or parameter transformations. Scenarios simulate real-world challenges, such as incomplete user inputs or conflicting requirements.
Examples of complex scenarios from medium and hard questions are included in Appendix E.
@@
The CallNavi dataset combines automation with human oversight, resulting in a benchmark that is both realistic and challenging. By spanning easy, medium, and hard tasks across diverse real-world domains, as outlined in Table 2, the dataset evaluates LLM capabilities in scenarios requiring robust task routing, contextual understanding, and API management. The dataset format and metadata details are provided in Appendix A.
Table 2: CallNavi dataset domains, questions and difficulties statistics table.
{tblr}
cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=3c, vlines, hline1,3-14 = -, hline2 = 4-6, Domain & API
Functions Questions Difficulty
Easy Medium Hard
Banking 91 115 70 28 17
Shopping 81 65 41 17 7
Logistics 46 65 40 17 8
Aviation 48 80 44 24 12
Healthcare 20 47 31 10 6
Public Services 82 85 50 27 8
Human Resources 20 35 21 13 1
Hotel Industry 49 65 40 19 6
Insurance 42 60 40 11 9
Telecommunications 100 112 79 22 11
Overall 579 729 456 188 85
4.1 Dataset
The CallNavi dataset [^1] https://anonymous.4open.science/r/CallNavi-783E evaluates LLMs’ task routing and API calling capabilities across multiple domains. As shown in Table 2, it contains 729 questions of varying difficulty and API interaction complexity, along with 579 distinct API functions. Questions are categorized into easy, medium, and hard levels:
  • Easy (456 questions): Require a single API call to fulfill the task.
  • Medium (188 questions): Involve multiple APIs within the same question, with all necessary parameters provided in the context.
5 Metrics and Evaluation
5.1 API Parameters AST Match
Our study utilizes Abstract Syntax Tree (AST) evaluation to assess models’ ability to generate accurate JSON outputs for API calls. The format of the output JSON and parameters follows a structure similar to the BFCL dataset in Section A Yan et al. (2024). We parse the generated JSON string into an object and compare each component with the ground truth, such as the API list and parameters.
  • AST Exact Match: Whether the entire parsed object, including its structure and content(values), is identical to the ground truth.
Figure 2: Example of evaluation pipeline
This multi-step evaluation ensures a thorough assessment of both the accuracy of API function calls and the structural integrity of the parameters, allowing for a granular analysis of the model’s performance.
Figure 3: Pipeline of AST Match Score
5.2 LLM-as-a-Judge Evaluation
We also use GPT-4o language models to evaluate whether the generated JSON outputs correspond accurately to the ground truth Zheng et al. (2023). This approach aims to observe if LLMs can perform such evaluation tasks with high precision. By leveraging an LLM for this purpose, we assess its capability to compare and validate structured data, thereby determining its effectiveness in automating the evaluation process. An example prompt is shown in appendix B.2.
5.3 Stability Score
  • $N$: Total number of outputs (samples).
The stability score is calculated as:
Stability $\text{Stability Score}=\frac{F_{1}-F_{2}}{N-F_{2}}$
To ensure reliable comparisons and reduce errors, we preprocess the outputs by removing unnecessary spaces, newlines, and formatting inconsistencies, converting the text to lowercase, and stripping extraneous characters that could cause mismatches. More detailed analysis can be found in the appendix D.
While ‘freq@topk‘ is often used to evaluate model performance, it fails to capture stability in the generated outputs. For example, if a model produces the sequence "AABBC" across multiple runs, ‘freq@topk‘ might assign a high score of 0.4 because the most frequent token ("A") appears 40% of the time. However, this sequence is not stable, as no single output consistently dominates. In contrast, our stability score focuses on the dominance of the most frequent output, offering a better measure of a model’s reliability in structured tasks.
To further evaluate the stability of the model’s outputs, we calculate the average Levenshtein distance between the first answer and each subsequent outputZhang et al. (2017). We normalize Levenshtein distance using the following formula.
6 Experiments and Benchmark
6.1 Models
To evaluate the benchmark, we selected models based on their performance, architecture, and relevance to function-calling tasks. The selection criteria focused on general-purpose and fine-tuned models optimized for function calling or JSON generation, ensuring a well-rounded comparison between zero-shot and fine-tuned capabilities. Models like BART and traditional retrieval-based approaches were excluded as they lack the ability to select APIs from extensive lists or structure API calls sequentially, which are critical for the complexity of this task.
Table 3 organizes the selected models into four groups: commercial models, middle-large models (10B+ parameters), small models (5B–10B parameters), and lightweight models (under 5B parameters).
Model Name Origin Parameters
GPT-4o OpenAI Achiam et al. (2023) N/A
GPT-4o-mini OpenAI Achiam et al. (2023) N/A
Gemini 1.5 Flash Google Reid et al. (2024) N/A
LLaMA 3.1 Meta AI Meta-AI (2024a) 70B
Command-R Command AI Cohere 35B
Gemma2 Google Team et al. (2024) 27B
Mistral-Small Mistral AI Ai (2024a) 22B
Phi3 Microsoft Abdin et al. (2024) 14B
Mistral-Nemo Mistral AI Ai (2024b) 12B
Gemma2 Google Team et al. (2024) 9B
LLaMA 3.1 Meta AI Meta-AI (2024a) 8B
xLAM Salesforce Zhang et al. (2024) 7B
NemoTron-Mini NVIDIA Paul (2024) 4B
Phi3.5 Microsoft Trufinescu (2024) 3B
LLaMA 3.2 Meta AI Meta-AI (2024b) 3B
NexusRaven Nexusflow.ai Nexusflow.ai (2023) 13B
Gorilla Berkeley Patil et al. (2023) 7B
Table 3: Comparison of General Purpose LLMs.
We choose 2 fine-tuned Function Calling models for testing which have top performance on the BFCL leaderboard: NexusRaven and Gorilla OpenFunctions v2. Then we found that Firefunction v2 on the leaderboard cannot accept that long API calling list in the pilot test Garbacki and Chen (2024).
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=4, cell17 = r=2, cell18 = r=2, cell19 = c=5, cell114 = r=2, cell31 = r=3, cell61 = r=6, cell121 = r=3, cell151 = r=3, cell181 = r=2, cell201 = r=3, hline1,23 = -0.08em, hline2 = 3-6,9-13, hline3 = -, Category & Models API Calling Routing
Exact Match Syntax
Validity Structural
Accuracy API Calling with Parameters
AST Match Overall
GPT Score
Easy Medium Hard All Easy Medium Hard All Avg. Macro Avg.
Commercial
Large
General
Middle
Light
Phi3.5 3B 0.723 0.340 0.188 0.562 0.004 0.0 0.0 0.0 0.0 0.0 0.0 0.002
Table 4: Benchmark Results Table, Fine-tuned model with star* means using Asynchronous Generation with GPT4o calling prediction. Macro Avg. means the arithmetic mean of 3 difficulties.
6.2 Environment
6.3 Pipeline
Our evaluation pipeline begins with the creation of prompts based on two templates, as detailed in the appendix. The first template in the appendix B.1 focuses on retrieving the API calling list, which corresponds to the "API" list in the ground truth. This prompt instructs the model to identify which API calls should be used and in what order. The second template is designed to generate the full API calling JSON, including the parameters for each call.
6.4 Benchmarks Results
The results presented in Table 6.1 highlight the performance of various models across different aspects of API function calling, including API calling routing accuracy, syntax validity, structural accuracy, and API parameter match through AST evaluation. OpenAI’s models, GPT4o and GPT4o mini, consistently outperform the others, particularly in syntax validity (0.993 and 0.994, respectively) and overall GPT score (0.913 and 0.908). Both models also demonstrate strong structural accuracy and API parameter AST match, especially in easier tasks. And Gemini 1.5 Flash follow closely in these metrics.
Our analysis demonstrates that the Pearson correlation between the "GPT Score" and the "All Avg." column in the "Parameter AST Match" section is 0.934, with a p-value of 4.40e-08. This indicates a very strong positive correlation, suggesting that higher GPT scores are closely associated with better average AST match performance. The results in our table closely align with those from the Berkeley Function Calling Leaderboard [^2] https://gorilla.cs.berkeley.edu/leaderboard.html, which assesses LLMs’ performance in API or function-calling tasks. In both evaluations, models like OpenAI’s GPT4o stand out for their high syntax validity and overall accuracy, as reflected in our table where GPT4o scores above 0.9 in both categories. This matches the leaderboard’s top models, which also excel in AST evaluations and execution accuracy. In contrast, fine-tuned models such as FireFunction V2 in our results show weaker performance in API calling accuracy and AST matching, a trend similarly observed with fine-tuned models like Gorilla OpenFunctions on the Berkeley leaderboard, particularly in more complex API scenarios. Both evaluations emphasize the challenges faced by fine-tuned models in handling complex function scenarios or multi-step API calls, highlighting the need for improvement in these areas.
6.4.1 Stability Test
In our stability experiments, we ran 5 times referring to the previous study Tian et al. (2023), and the stability results are in Table 7 in Appendix. By these metrics results, we obtain a numerical comparison that reflects the stability differences between outputs.
As mentioned in Section 5.3, a higher Election Stability Score indicates greater absolute consistency in the model’s outputs across multiple runs. A high Levenshtein Stability Score means similar between the same input in text generation output. Commercial models also perform better in Table 7 in Appendix.
7 Zero Shot Improvement
7.1 Calling + Parameters Asynchronous Generation
To address these challenges, we propose combining the strengths of general-purpose LLMs and fine-tuned models. Specifically, a general LLM selects the relevant APIs based on the input prompt, leveraging its superior understanding in identifying appropriate API calls. These selected APIs are then provided to a fine-tuned/LAM model, which focuses on generating the correct API calls along with the necessary parameters. This sequential process allows the general LLM to efficiently handle API selection, while the fine-tuned model concentrates on accurately producing API calls and parameters within a more manageable context.
As demonstrated in Table 5 (models marked with an asterisk*), this combined approach with GPT-4o routing significantly improves performance. Separating the tasks of API selection and parameter generation enhances the models’ ability to handle complex API calling tasks more effectively.
Table 5: Asynchronous Generation results for LLMs
Models easy medium hard overall
Gorilla v27B 0.682 0.005 0.000 0.427
xLAM-fc-7B 0.714 0.462 0.188 0.588
Gemma2 0.723 0.457 0.164 0.589
llama3.1 0.714 0.457 0.164 0.584
mistral-small 0.728 0.436 0.294 0.602
mistral-nemo 0.712 0.308 0.141 0.541
phi3:14b 0.019 0.005 0.011 0.015
command-r 0.633 0.547 0.223 0.563
llama3.2 0.462 0.297 0.082 0.375
nemotron-mini 0.208 0.01 0.000 0.133
LLM w/ itself as as router Gemma2:27b 0.598 0.59 0.341 0.566
Mistral-small 0.684 0.382 0.235 0.554
Command-r 0.621 0.505 0.247 0.547
7.2 Backward Inference Thinking
To optimize the API selection and calling process, we implemented an approach inspired by CauseJudgerHe and Lu (2024) and Reverse ChainZhang et al. (2023) on the given task. The model is then guided to confirm whether this final API satisfies all the requirements in the question and whether all necessary information is included in the parameters. If the model identifies missing information, it is prompted to think through which additional API calls could provide the missing data. This iterative process continues until all information required by the question is gathered and all requirements are met. By breaking the task into these smaller, more manageable steps, the LLM can effectively plan and execute a series of API calls that better address the complexity of the scenario. This method has led to a performance improvement of approximately 30% in hard-level questions API callings in Table 6 test with GPT-4o/mini, where the challenges of multiple selections and nested calls are more prevalent. This optimization enhances the model’s ability to handle intricate, real-world API calling tasks more effectively.
Table 6: Backward Thinking performance in High difficulty calling in GPT-4o and GPT-4o-mini.
Model API Calling Routing API Calling with Parameters
Original
Backward
Thinking
Original
Backward
Thinking
GPT4o 0.611 0.894 0.388 0.729
GPT4o mini 0.564 0.847 0.364 0.482
8 Discussion and Conclusion
Our dataset introduces significant challenges, particularly in medium and hard questions, where models must accurately select APIs from a large candidate pool and generate parameters in multi-step and nested contexts. This complexity, absent in simpler datasets, reveals the limitations of fine-tuned models traditionally trained on smaller API sets. The higher variability and complexity of our data underscore the need for more diverse and robust training paradigms.
In evaluating models on API calling routing and parameter JSON generation, we observed distinct behaviors. GPT-4o consistently excelled in both tasks, reinforcing its utility for complex language tasks. Open-source models like LLaMA 3.1 and Gemma2 performed well in API routing but struggled with parameter generation, making them cost-effective options for simpler routing tasks. However, smaller models (<10B parameters) exhibited instability in complex tasks, often producing inconsistent or incomplete outputs, highlighting their limitations in high-accuracy, multi-step API scenarios.
Despite these advancements, none of the models, including GPT-4o, achieved perfect results in handling the hardest scenarios, indicating that fully mastering intricate API calling tasks remains an open challenge.
Conclusion
This work addresses the challenges of API function calling and parameter generation in LLMs by introducing a novel dataset of over 500 APIs and 700 questions, alongside a comprehensive benchmark for general-purpose and fine-tuned models. We also use two optimization strategies asynchronous generation and backward inference to enhance multi-step API selection and parameter generation. Our findings reveal current limitations and open avenues for advancing LLM-driven function calling in theory and practice.
Limitations
While our work presents important advancements in API function calling routing and parameters generation for LLMs, there are several limitations to acknowledge. First, the scenarios we used in our dataset are simplified versions of real-world situations and may not fully capture the complexities that arise in actual applications. This could limit the practical applicability of our results in highly dynamic or unpredictable environments. Additionally, our team has limited expertise in designing datasets across the many diverse domains included in this study, which may have led to oversights in representing certain domain-specific challenges accurately.
Another limitation is related to the performance of the models we tested. While we selected well-regarded LLMs for benchmarking, they may not have performed at their best due to factors such as suboptimal configurations or fine-tuning for different tasks. Lastly, we acknowledge the absence of a mock server in our experimental setup, which could have enhanced the realism of the API calling process. Incorporating a mock server to simulate real-time API interactions would provide a more accurate assessment of how these models handle real-world API calling tasks. Future iterations of this work should address these limitations to provide more comprehensive and applicable insights.
Ethic Statment
In the development of our dataset, we ensured that no private or sensitive information was used, and all data was mock information, generated in compliance with GDPR. This guarantees that no data leakage or privacy concerns are associated with our work. Additionally, we made efforts to minimize the environmental impact of our research by optimizing model training and evaluation processes, thus reducing the carbon footprint associated with our benchmarking tasks. Where possible, we utilized energy-efficient infrastructure to further support this initiative. Our commitment to ethical practices extends across all aspects of this research, ensuring data privacy and environmental responsibility.
Appendix A Meta Data
The first part is the API name, description, and parameters.

### v2

# CallNavi, A Challenge and Empirical Study on LLM Function Calling and Routing

@@
Abstract.
API-driven chatbot systems are increasingly integral to software engineering applications, yet their effectiveness hinges on accurately generating and executing API calls. This is particularly challenging in scenarios requiring multi-step interactions with complex parameterization and nested API dependencies. Addressing these challenges, this work contributes to the evaluation and assessment of AI-based software development through three key advancements: (1) the introduction of a novel dataset specifically designed for benchmarking API function selection, parameter generation, and nested API execution; (2) an empirical evaluation of state-of-the-art language models, analyzing their performance across varying task complexities in API function generation and parameter accuracy; and (3) a hybrid approach to API routing, combining general-purpose large language models for API selection with fine-tuned models and prompt engineering for parameter generation. These innovations significantly improve API execution in chatbot systems, offering practical methodologies for enhancing software design, testing, and operational workflows in real-world software engineering contexts.
Function Calling, Large Language Models, Chatbot, Benchmark
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Türkiyeisbn: 978-1-4503-XXXX-X/2025/XXccs: Software and its engineering Empirical software validationccs: Computing methodologies Information extractionccs: Software and its engineering Software design engineering
1. Introduction
@@
Modern conversational AI systems, such as chatbots, rely on accurate API calling to enable effective user interactions, as shown in Figure 1. Beyond generating simple API calls, models must handle complex scenarios involving selecting the correct API from extensive lists, orchestrating multiple sequential calls, and managing nested API interactions. While progress has been made in generating syntactically correct single API calls, there is limited focus on generating sequences of API calls with logical dependencies in long description context, a crucial requirement for real-world applications.
@@
Large Language Models (LLMs), such as GPT-4 (Achiam et al., 2023) and Llama (Touvron et al., 2023), have demonstrated impressive capabilities in various natural language processing tasks. These models excel at generating coherent and contextually relevant responses, but their ability to produce structured outputs, such as API calls, program code, or other machine-readable formats, remains a challenging frontier. Structured output generation requires adherence to predefined syntactic and semantic rules, making it more constrained than generating free-form text(Liu et al., 2024b).
@@
Recent advancements have explored structured output generation in applications such as code generation (Chen et al., 2021), table completion (Herzig et al., 2020), and multi-turn dialogue (Budzianowski et al., 2018). Tools like CodeX (Chen et al., 2021) and AlphaCode (Li et al., 2022) focus on generating functionally valid code, while methods like chain-of-thought prompting (Wei et al., 2022) and tool-augmented reasoning frameworks (Qin et al., 2023) enhance reasoning in complex tasks. These methods highlight the potential of LLMs for tasks requiring step-by-step reasoning and structured output generation.
@@
Early studies, such as BotBase (Zamanirad et al., 2017), explored translating natural language into API calls, laying the groundwork for automating tool use. More recent benchmarks, including API-Bank (Li et al., 2023), ToolEyes (Ye et al., 2024), and BFCL (Yan et al., 2024), evaluate LLMs on API execution. However, these datasets often operate with small API candidate pools or lack scenarios involving nested or interdependent API calls. For example, API-Bank assesses tool-augmented models but limits API candidates to fewer than five per task. Similarly, ToolBench (Liu et al., 2024a) and ToolEyes evaluate multi-tool scenarios but do not support tasks requiring highly interdependent API calls.
@@
To address these gaps, we propose CallNavi, a novel benchmark designed to evaluate LLMs on:
  • Selecting APIs from an unfiltered list of over 100 candidates;
  • Executing multiple sequential API calls;
  • Handling nested API interactions.
@@
CallNavi introduces real-world complexity by simulating unfiltered API selection and combining generation with routing tasks in a long context input. It categorizes questions into easy, medium, and hard levels, enabling a granular evaluation of model capabilities across varying task complexities. Additionally, we propose new metrics, including a stability score, to measure prediction consistency across multiple runs.
@@
We benchmark 18 LLMs, encompassing commercial, general-purpose, and fine-tuned models, on CallNavi. Our findings provide insights into the strengths and limitations of current models, laying a foundation for advancing API selection and function calling.
Figure 1. Example of API Calling pipeline via LLM
2. Related Work
Generating and executing accurate API calls is crucial to integrating LLM into real-world conversation applications. Existing benchmarks, such as API-Bank (Li et al., 2023), ToolEyes (Ye et al., 2024), and ToolBench (Liu et al., 2024a), evaluate API selection and execution capabilities but often rely on prefiltered API candidate pools, lack nested API tasks, or focus on narrow domain coverage. In contrast, CallNavi introduces unfiltered API selection with over 100 candidates, multi-call tasks, and nested API scenarios across 10 diverse domains. Table 1 highlights these distinctions, demonstrating how CallNavi addresses limitations in existing benchmarks by introducing realistic complexity and structured difficulty levels.
Structured output generation, a critical capability for API function calling, has seen significant advancements. MetaGPT (Hong et al., 2023) and CodeAgent (Tang et al., 2024) emphasize task decomposition and multi-step reasoning, improving performance in complex workflows. Techniques like constrained generation (Beurer-Kellner et al., 2024) and grammar-aware Seq2Seq models (Dong et al., 2023) improve structured output reliability, aligning with CallNavi’s focus on evaluating structured reasoning and accuracy.
Stability in LLM predictions is another vital area of research. Although traditional metrics such as freq@topk assess prediction reliability, they fail to capture consistency across multiple runs fully. Inspired by prior stability-focused studies (Du et al., 2023; Tian et al., 2023), CallNavi introduces a stability score to quantify prediction consistency, complementing traditional metrics like AST match and exact match.
From a software engineering perspective, function-calling tasks align with modularity and abstraction principles, emphasizing decomposition into manageable sub-tasks. Early works, such as BotBase (Zamanirad et al., 2017), synthesized API calls from natural language, laying the groundwork for modern tools like Gorilla (Ji et al., 2024), ToolLLM (Qin et al., 2023), and ToolAlpaca (Tang et al., 2023). Recent efforts like StableToolBench (Guo et al., 2024) and $\tau$-bench (Yao et al., 2024) highlight challenges in tool learning and real-world tool-agent-user interactions.
API recommendation systems, such as those explored by Peng et al. (Peng et al., 2022), provide insights into ranking and selecting APIs. These systems complement CallNavi, which emphasizes multi-step workflows requiring careful API selection. Similarly, abstract syntax networks (Rabinovich et al., 2017) and benchmarks like BigCodeBench (Zhuo et al., 2024) advance structured code generation and semantic parsing, aligning with CallNavi’s emphasis on reasoning and logical consistency in nested tasks.
In summary, while prior work has laid a strong foundation for API function calling, CallNavi advances the field by addressing critical gaps such as unfiltered API selection, nested tasks, and stability evaluation. These contributions provide a robust framework for benchmarking LLMs in realistic and complex scenarios.
@@ 2. Related Work
Table 1. Comparison of CallNavi with existing API function-calling benchmarks test set.
@@
3. Research Questions
@@ 4. CallNavi Dataset
All generated API functions were manually reviewed for accuracy, consistency, and relevance. i.e.:
@@
  • Coverage of straightforward, single API usage scenarios.
@@
  • Hard Questions: Designed to address edge cases, ambiguous queries, and nested API calls requiring advanced reasoning. Scenarios simulate real-world challenges, such as incomplete user inputs or conflicting requirements.
@@
The CallNavi dataset combines automation with human oversight, resulting in a benchmark that is both realistic and challenging. By spanning easy, medium, and hard tasks across diverse real-world domains, as outlined in Table 2, the dataset evaluates LLM capabilities in scenarios requiring robust task routing, contextual understanding, and API management.
The first part of the metadata is a long JSON file with the API name, description, and parameters in the following format.